OpenStudy (anonymous):

Line GI bisects angle DGH so that the measurment of angle DGI is x-3 and the measurement of angle IGH is 2x-13. Find the value of x. PLEASE EXPLAIN THE ANSWER

OpenStudy (campbell_st):

since GI bisects that means the angle is cut into 2 equal angles or DGI = IGH so x - 3 = 2x - 13 solve for x

OpenStudy (anonymous):

I don't know how to solve for x though..... That's what I need help with.

OpenStudy (anonymous):

Wait, would x equal 10? :3

OpenStudy (anonymous):

 they are complementary angle so.. (x-3) + (2x-3) = 90 3x= 90+6 X = 32